OpenStudy (anonymous):

Evaluate tan^2 60 degrees + sin^2 45 degrees without using trigonometric tables or calculator. The answer is a. 5 b. 7/3 c. 1/2 d. 7/2

OpenStudy (anonymous):

See we have some standard values that we have to remember: Here I tell you: \[\large \sin(45) = \frac{1}{\sqrt{2}}\] \[\large \tan(60) = \sqrt{3}\] So square each of them and add and tell me what do you get ?

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Absolutely..

OpenStudy (anonymous):

Did not you get 3.5 ??

OpenStudy (anonymous):

Yes, is that the correct answer?

OpenStudy (anonymous):

D. Thank you!

OpenStudy (anonymous):

In choices divide the fractions : 7/3 = ??? 7/2 = ???

OpenStudy (anonymous):

Yes you are right...